OpenStudy (anonymous):

Can you guys explain poles, roots and horizontal asymptotes and how to determine them in an equation/graph?

OpenStudy (anonymous):

Roots are anywhere such that: \[ f(x)=0 \]

OpenStudy (anonymous):

Poles are anywhere such that: \[ \left|\lim_{x\to a}f(x) \right|= \infty \]

OpenStudy (djcall11):

i cant really explain it that well but i have a question are you work on something in flvs.com and maybe this video can help you https://www.youtube.com/watch?v=K2Bku1DsIto

OpenStudy (anonymous):

Horizontal asymptotes are anywhere such that: \[ \lim_{|x|\to\infty} f(x) = L \]

OpenStudy (perl):

poles occur when the denominator is equal to zero, in the simplified or canceled version of the rational function

OpenStudy (anonymous):

are poles vertical asymptotes?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so say you have a rational expression...

OpenStudy (anonymous):

For a rational expression \[ f(x)=\frac{g(x)}{h(x)} \]The roots of \(h(x)\) (\(h(x)=0\)), are the poles of \(f(x)\).

OpenStudy (anonymous):

I don't know something like \[4((x+2)(x-4))/(x-2)^2\]

OpenStudy (anonymous):

The exception being if \(g(x)\) also has a root at the same place.

OpenStudy (perl):

you have to be careful to cancel common factors in the numerator and denominator, or you get a pseudo pole (or just a hole)

OpenStudy (anonymous):

Since \(h(2)=0\) and \(g(2) = -16\neq 0\), then \(f(x)\) has a pole at \(2\).

OpenStudy (anonymous):

The limit, from both the left and right, approaches \(-\infty\).

OpenStudy (anonymous):

The roots of \(f(x)\) are the same as the roots of \(g(x)\). So in this case they are \(-2\) and \(4\).

OpenStudy (anonymous):

For horizontal asymptotes, it is a bit more tricky, I suppose.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

First you want to know the degree of the polynomials. \(g(x)\) is degree 2, and \(h(x)\) is degree 2.

OpenStudy (anonymous):

You'll only need to look at the coefficient of the highest degree.

OpenStudy (anonymous):

if they aren't of the same degree then there isn't a horizontal asymptote?

OpenStudy (anonymous):

Cancel out all but the highest degree in each factor:\[ \frac{4((x+2)(x-4))}{(x-2)^2} \to \frac{4((x+\cancel{2})(x-\cancel{4}))}{(x-\cancel{2})^2} = \frac{4(x)(x)}{(x)^2} = \frac{4x^2}{x^2}=4 \]

OpenStudy (anonymous):

This only really works for rational functions.

OpenStudy (anonymous):

You'll end up with something of the form\[ \frac{ax^m}{bx^n} \]

OpenStudy (anonymous):

When \(m=n\), then your horizontal asymptote is at \(a/b\). When \(m>n\), then you don't have a horizontal asymptote. When \(m<n\), then your horizontal asymptote is at \(0\).

OpenStudy (anonymous):

okay so poles are when the denom = 0, roots are where the num = 0 and hoz asymptotes are when the degree in the num is less than or greater than the denom.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

for rational expressions anyways

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

For a graph...

OpenStudy (anonymous):



OpenStudy (anonymous):